There are 7 more boys than girls in a class of 25.how many girls are in the class?


1
Expert's answer
2021-08-26T09:02:51-0400

Solution.

Let x be the number of girls in a class of 25.

Then there are x+7 boys in a class of 25.

We will have equation

x+x+7=25.

Solve this equation


"x+x+7=25\newline
2x+7=25\newline
2x=18\newline
x=9"

Thus, there are 9 girls and 9+7=16 boys in a class of 25.

Answer. 9 girls.